OpenStudy (anonymous):

Simplify and determine the coefficient of (-2/3 x)(3y)(-2x). -4 1/3 1 4

OpenStudy (anonymous):

4?

OpenStudy (haleyelizabeth2017):

-2/3*-2=?

OpenStudy (here_to_help15):

you are correct :D

OpenStudy (anonymous):

oh so its 1/3?

OpenStudy (here_to_help15):

Its 4

OpenStudy (haleyelizabeth2017):

No....I am just explaining how to find it lol

OpenStudy (here_to_help15):

lol

OpenStudy (haleyelizabeth2017):

@Here_to_Help15, would you care to explain?

OpenStudy (anonymous):

oh haha ok thnx:))

OpenStudy (here_to_help15):

Gulp* of course haha ;)

OpenStudy (here_to_help15):

you just simplify it to (-2/3x)(-6xy)

OpenStudy (haleyelizabeth2017):

I did it differently....

OpenStudy (here_to_help15):

We are 2 different people hun

OpenStudy (here_to_help15):

4xy

OpenStudy (haleyelizabeth2017):

(-2/3x)(-2)=(4/3x)*(3y) \(4\)xy Well no der

OpenStudy (here_to_help15):

lol

OpenStudy (haleyelizabeth2017):

Whoops I missed the x....it would be \((4/3)x^2\), not (4/3)x

OpenStudy (here_to_help15):

ya :)

OpenStudy (anonymous):

wait what??

OpenStudy (haleyelizabeth2017):

I'm so lost.....

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (haleyelizabeth2017):

I think it is \(4x^2y\) so it is still 4.....

OpenStudy (anonymous):

ok thnx

OpenStudy (haleyelizabeth2017):

You are welcome :)